rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#1841 — FS#5814 — peering orange th2

Attached to Project— Network
Incident
Whole Network
CLOSED
100%
Several customers raise the problem of
connection to Orange. This night we have
carried out the work of migration of th1 to
th2 and one of the two peering of Paris was
transferred to th2.
http://status.ovh.net/?do=details&id=1832

We look for the origin of the problem,it should not be a long way off
According to the graphs,the flow in a channel port
of 2x10G is not equal between the two 10G ports. We cut the port which we think is at the origin of the problem.

The overall flow port channel increased by 30%.
99% it is the port that poses the problem

If you do not have , or if you still have
the problem,would you send a confirmation email to noc@ovh.net, oles@ovh.net
thank you in advance!

Date:  Saturday, 01 October 2011, 20:47PM
Reason for closing:  Done
Comment by OVH - Friday, 23 September 2011, 16:07PM

A customer confirms the problem resolution.
Thank you for the feedback.

We will check why this port is not functioning.


Comment by OVH - Monday, 26 September 2011, 17:23PM

We changed the xenpack on the port 9/3. We try to reactivate the port.


Comment by OVH - Monday, 26 September 2011, 17:24PM

Same problem. We're looking if it's possible to improve the
levels of attenutation on the link.


Comment by OVH - Monday, 26 September 2011, 17:29PM

We checked another default on the optical link of the first port.
The 2 ports are currently down. The traffic is flowing via gsw.


Comment by OVH - Monday, 26 September 2011, 23:00PM

We checked totally the internal optical cabling for these two links, and replaced several garters.
We didn't note problems anymore. The two ports are properly balanced concerning traffic.
If you detect new inconveniences to Orange,please contact directly noc@ovh.net.


Comment by OVH - Monday, 26 September 2011, 23:03PM

Orange remounts us errors from their side on one of the two ports. We are forced to cut it at this time to avoid any impact on traffic. A technician will intervene tomorrow morning in order to fix the problem.


Comment by OVH - Tuesday, 27 September 2011, 16:38PM

We intervene on the physical interco.


Comment by OVH - Saturday, 01 October 2011, 20:47PM

The problem is fixed.